OpenStudy (anonymous):

solve the equation...x(x-4)=12 then x(x-4)-12=0...wha comes after this

OpenStudy (anonymous):

x^2-4x-12=0 x^2-6x+2x-12=0 ........factorize...

OpenStudy (anonymous):

x=-2 x=6

OpenStudy (anonymous):

yup whatever

OpenStudy (anonymous):

x^2-4x-12=0 x^2+(2-6)*x+(2*(-6))=0 (x+2)*(x-6)=0